OpenStudy (anonymous):

Help please! Doing Pr-Cal Eliminate the parameter. x = square root of t y = 3t + 7

OpenStudy (campbell_st):

ok... so if \[x = \sqrt{t}\] square both sides of the equation \[x^2 = t\] now substitute this into y = 3t + 7

OpenStudy (anonymous):

x = sqrt(t) implies t = x^2 so y = 3x^2 + 7. Notice that x is non-negative, so you have only half of a parabola

OpenStudy (anonymous):

Thanks everyone! That makes more sense!